### v1

# Unified Lorentz-covariant Poisson Bracket for the Electrodynamics of a Point Particle

Abstract
Using the multisymplectic Hamiltonian formalism, we propose a Poisson bracket for the electromagnetic field that, in addition to satisfying the restricted principle of relativity, reproduces well-established results from the standard formulations of field theories. To this end, it is necessary to translate the multisymplectic description of the system to the field’s momentum representation, where a bilinear and Lorentz-covariant form emerges in a straightforward manner. Consequently, we establish a description for a system consisting of an electromagnetic field and a point particle interacting with each other that is based on a single Lorentz-covariant Poisson bracket with respect to all the canonical variables of the system. The results obtained from this proposal offer a solid foundation for the development of a Lorentz-covariant quantization process.
Keywords: Field theory, classical electrodynamics, multisymplectic Hamiltonian formalism, Lorentz-covariant Poisson bracket
1 Introduction
Since the early objections raised by Born [1], the incompatibility between the requirement of full consistency with the restricted principle of relativity and the use of canonical quantization in modern field theories has been addressed on numerous occasions without yet reaching a definitive solution. The crux of the matter lies in that canonical quantization requires a Hamiltonian (or canonical) formulation of the theories and, as a consequence, implies a primacy of time as fundamental independent variable. This incompatibility is only apparent, naturally, since quantum field theories do describe relativistic phenomena. However, the covariance of such theories must be verified at each stage of their development by means of generally laborious demonstrations.
Constructing a Lorentz-covariant bilinear form that generalizes the Poisson bracket (pb) of classical mechanics when dealing with fields is another open question in theoretical physics that is closely related to the previous issue. In this regard, the studies done by Peierls [2] and by Marsden et al. [3] are of great significance, as they clearly illustrate that there is no unique or straightforward way to construct a pb with the aforementioned characteristics. Multiple proposals for Lorentz-covariant pbs have been published over the years for both classical and quantum fields (see, for example, [2, 3, 6, 7, 8, 9, 10, 4, 5] and references therein), indicating that no general consensus exists on this matter.
Hence, generalizing the Hamiltonian formalism in such a way that it can underpin a manifestly Lorentz-covariant quantization process is a topic of major interest. A strong candidate for this generalization is the multisymplectic Hamiltonian formalism (mhf), whose main feature is to manifestly preserve the covariance of the canonical description of fields through the introduction of a generalized momentum per field derivative. The seminal papers on the mhf were published by de Donder [11] and by Weyl [12] in the 1930s. However, after nearly a century of existence, the use of this formalism to construct field theories is not common practice, since its development beyond the basic definitions is not straightforward and still presents formal lacunae [13]. In particular, the mhf has as an important drawback that it allows to define more than one pb [13, 14, 15, 9, 16, 17]. Nonetheless, it has recently come to our attention that it is possible to construct a Lorentz-covariant pb in a straightforward manner for the electromagnetic field by translating the multisymplectic description of the field to its momentum representation [18]. Exploring this possibility is the main subject of the present work.
The paper is organized as follows. In Sect. 2, after briefly reviewing the fundamentals of describing a system composed of an electromagnetic field and a single point particle using the mhf, we translate the generated description to the momentum representation. The central part of the paper is contained in Sect. 3 and deals with the proposal of a pb with respect to the canonical variables of the field that is capable not only of satisfying the restricted principle of relativity but also of reproducing well-established results from the conventional formulations of field theories. Subsequently, in Sect. 4, the non-covariant form of the proposal is analyzed and compared with the non-covariant pb that is associated with the electromagnetic field. Lastly, we outline a description of the complete system on the basis of the found pb and briefly discuss its main implications in Sect. 5.
2 Multisymplectic Hamiltonian Description
We start by properly setting up the system to be investigated as an electromagnetic field and a single point particle interacting with each other in a four-dimensional Minkowski space-time, with metric tensor $\eta_{\mu\nu}$ and signature (+,,,)(+,-,-,-). The action of such a system can be stated as
$S=\int{d}^{4}x\hskip 1.42262pt(\mathcal{L}_{\mathrm{field}}+\mathcal{L}_{\mathrm{part}}+\mathcal{L}_{\mathrm{int}}),$ (1)
where $\mathcal{L}_{\mathrm{field}}$, $\mathcal{L}_{\mathrm{part}}$, and $\mathcal{L}_{\mathrm{int}}$ are the Lagrangian densities associated with the free propagation of the field, the free motion of the particle, and the particle++field interaction, respectively. Thus, we have that
$\mathcal{L}_{\mathrm{field}}=-\frac{1}{8\pi{c}}\partial_{\mu}A_{\nu}\partial^{\mu}A^{\nu},$ (2)
$\mathcal{L}_{\mathrm{part}}=-\frac{m_{0}c^{2}}{2}\int{d}\tau\hskip 1.42262pt\delta^{(4)}(x_{\alpha}-u_{\alpha}),$ (3)
and
$\mathcal{L}_{\mathrm{int}}=-\frac{e}{c}A_{\mu}\int{d}\tau\hskip 1.42262pt\dot{u}^{\mu}\delta^{(4)}(x_{\alpha}-u_{\alpha}),$ (4)
with $A_{\mu}$ the field’s vector potential and $m_{0}$, $e$, $u_{\mu}$, and $\tau$ the rest mass, electric charge, position, and proper time of the particle, respectively [^2] The Lagrangian density in Eq. (2) does not correspond to the gauge-invariant form commonly used in electrodynamics that is given in terms of the electromagnetic tensor. However, both Lagrangian densities differ only by a quantity that reduces to zero when integrated over the whole space-time and when considering the Lorentz condition, $\partial_{\mu}A^{\mu}=0$, so that the selection of one option or the other does not modify the action of the system..
Thereupon, the mhf implies a generalized conjugate momentum for $A_{\mu}$ given by
$\theta_{\mu\nu}=-\frac{1}{4\pi{c}}\partial_{\mu}A_{\nu},$ (5)
while the application of the usual Hamiltonian formalism results for the mechanical subsystem in the well-known canonical momentum
$p_{\mu}=-m_{0}\dot{u}_{\mu}-\frac{e}{c}A_{\mu}(u_{\alpha}).$ (6)
Consequently, we obtain that the de Donder-Weyl function, i.e., the Lorentz-covariant Hamiltonian density, for the system is
$\mathcal{H}=-2\pi{c}\hskip 0.7113pt\theta_{\mu\nu}\theta^{\mu\nu}-\frac{1}{2m_{0}}\int{d}\tau\left[p_{\mu}+\frac{eA_{\mu}}{c}\right]\left[p^{\mu}+\frac{eA^{\mu}}{c}\right]\delta^{(4)}(x_{\alpha}-u_{\alpha}),$ (7)
which leads to two pairs of Lorentz-covariant first-order Hamilton’s equations: one pair for the field’s variables and the other for those of the point particle. The set of these equations consistently reproduces the manifestly Lorentz-covariant forms of Maxwell’s equations and the particle’s equation of motion.
Now, another possible Lorentz-covariant description of the system can be constructed in terms of the normal modes of oscillation of the field by expressing the retarded solution of Maxwell’s equations as
$A_{\nu}=\frac{1}{8\pi^{3}}\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha}){\mathcal{A}}_{\nu}e^{-{i}k_{\mu}x^{\mu}}+\mathrm{c.c.},$ (8)
where the Dirac $\delta$-function $\delta(k_{\mu}k^{\mu})$ and the Heaviside step function $\Theta(k_{0})$ ensure that the dispersion relation $k_{0}=|\mathbf{k}|$ is satisfied, and ${\mathcal{A}}_{\nu}$ is a vector amplitude given by
${\mathcal{A}}_{\nu}={\mathcal{A}}^{\mathrm{free}}_{\nu}+4\pi{i}e\int{d}\tau\hskip 1.42262pt\dot{u}_{\nu}\Theta(x_{0}-u_{0})e^{ik_{\mu}u^{\mu}},$ (9)
with $\mathcal{A}_{\mu}^{\mathrm{free}}$ a vector amplitude depending solely on $k_{\mu}$. Having expressed the solution of the field equation in this manner, one can easily observe that the presence of the particle results for the field in a superposition of plane waves with constant amplitudes due to its free propagation and of waves with variable amplitudes whose origin can be shown to be the radiation reaction of the particle (for a detailed proof, see [19]).
Then, recalling that the description of the field provided by the mhf in the momentum representation has been shown to be equivalent to its Lagrangian counterpart [18], we have that Eq. (7) can be rewritten as
$\mathcal{H}=\frac{1}{8\pi^{3}}\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})\tilde{\mathcal{H}}$, (10)
where
$\tilde{\mathcal{H}}=-\frac{k_{\mu}k^{\mu}}{4\pi{c}k_{0}}{\mathcal{A}}_{\nu}^{\ast}{\mathcal{A}}^{\nu}+\delta_{k}\int{d}\tau\left[p_{\mu}\dot{u}^{\mu}+\frac{m_{0}c^{2}}{2}\right]\delta^{(4)}(x_{\alpha}-u_{\alpha})+\frac{e}{c}\int{d}\tau\left[\mathcal{A}_{\nu}\dot{u}^{\nu}e^{-ik_{\mu}u^{\mu}}+\mathrm{c.c.}\right]\delta(x_{0}-u_{0}),$ (11)
with $\delta_{k}$ a $\delta$-function such that $\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})\delta_{k}=8\pi^{3}$, is the image of the system’s de Donder-Weyl function in the field’s momentum representation. This last function has the same information about the system as $\mathcal{H}$ and therefore constitutes the basis for this alternative description. Furthermore, the structure of Eq. (11) reveals that the de Donder-Weyl function is a Lorentz scalar proportional to the rest mass of the complete system, i.e., particle++field [^3] The origin of the proposed interpretation becomes evident when recalling that the dispersion relation for a massive field is $k_{\mu}k^{\mu}=\mu^{2}$, with $\mu$ being the mass parameter associated with the field. For purposes of comparison, examples of de Donder-Weyl functions in the momentum representation for other fields may be seen in Ref. [18]. Naturally, this identification implies a null value for the de Donder-Weyl function’s part associated with the free propagation of the electromagnetic field since the dispersion relation states in such a case that $k_{\mu}k^{\mu}=k_{0}^{2}-|\mathbf{k}|^{2}=0$.
The image of the de Donder-Weyl function in the momentum representation can equivalently be expressed in terms of the field’s canonical variables ${q}_{\mu}$ and $\pi_{\mu\nu}$ [18], defined as
${q}_{\mu}=\frac{i(\mathcal{A}_{\mu}^{\ast}-\mathcal{A}_{\mu})}{\sqrt{8\pi{c}k_{0}}},\quad\pi_{\mu\nu}=\frac{k_{\mu}(\mathcal{A}_{\nu}^{\ast}+\mathcal{A}_{\nu})}{\sqrt{8\pi{c}k_{0}}},$ (12)
and whose Hamilton’s equations are:
$\partial_{\mu}{q}_{\nu}=\frac{\partial\tilde{\mathcal{H}}}{\partial\pi^{\mu\nu}},\quad\partial_{\mu}\pi^{\mu\nu}=-\frac{\partial\tilde{\mathcal{H}}}{\partial{q}_{\nu}},$ (13)
respectively. As a consequence, the set of equations constituted by Eqs. (13) and Hamilton’s equations for $u_{\mu}$ and $p_{\mu}$ completely governs the electrodynamics of the whole system.
3 Proposal for a Lorentz-covariant Poisson Bracket
In the Hamiltonian formulation of classical mechanics, the evolution equation of any phase-space function is stated in terms of its pb with the Hamiltonian function of the mechanical system concerned, with Hamilton’s equations being special cases of such an evolution equation for the canonical positions and momenta. In particular, when considering the Hamiltonian function itself, one immediately finds that its evolution equation reduces to an equality between its total and explicit time derivatives. So, if we claim that the above description of the system is canonical, we must verify whether the de Donder-Weyl function satisfies a similar mathematical relationship between its derivatives. To this end, we observe that it can be established straightforwardly from Eqs. (10) and (13) that
$\partial_{\mu}\mathcal{H}=\partial_{\mu}\mathcal{H}\hskip 1.42262pt\Big{|}_{\mathrm{explicit}}+\frac{1}{8\pi^{3}}\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})\left[\frac{\partial{\tilde{\mathcal{H}}}}{\partial{q}^{\nu}}\frac{\partial{\tilde{\mathcal{H}}}}{\partial\pi_{\mu\nu}}-\frac{\partial{\tilde{\mathcal{H}}}}{\partial{q}^{\nu}}\frac{\partial{\tilde{\mathcal{H}}}}{\partial\pi_{\mu\nu}}\right],$ (14)
so that we conclude that the de Donder-Weyl function indeed satisfies an equality such as the one mentioned.
The former statement brings pbs back to the center of our discussion. As mentioned in Sect. 1, the mhf accepts multiple ways of defining such geometric objects, yet the structure of Eq. (14) strongly suggests proposing the bilinear form
$\{A,B\}=\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})V_{\mu}\left[\frac{\partial{A}}{\partial{q}^{\nu}}\frac{\partial{B}}{\partial\pi_{\mu\nu}}-\frac{\partial{B}}{\partial{q}^{\nu}}\frac{\partial{A}}{\partial\pi_{\mu\nu}}\right],$ (15)
where $V_{\mu}$ is to be determined, as the equal-time pb of functions $A$ and $B$ with respect to the canonical variables of the field [^4] A preliminary version of this proposal can be found in Ref. [18].. The introduction of $V_{\mu}$ is strictly necessary for $\{A,B\}$ to be antisymmetric and satisfy Leibniz’s rule and Jacobi’s identity exactly as a pb is required to do (see Refs. [3, 6, 7, 8, 9] for proposals with a structure similar to the one presented here). That being the case, applying Eq. (15) implies that the equal-time pb of $q_{\mu}$ and $\pi_{\mu\nu}$ is given by
$\{q_{\mu}(x_{0},k_{\alpha}),\pi_{\lambda\nu}(x_{0},k_{\alpha}^{\prime})\}=\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})V_{\lambda}\eta_{\mu\nu}\delta^{(4)}(k_{\alpha}-k_{\alpha}^{\prime}),$ (16)
from which two conclusions can be drawn. First, and as one could expect, the canonical variables associated with different modes of oscillation of the field are independent of each other. Second, we can rewrite Eq. (16) so that it resembles the canonical pb of non-relativistic classical mechanics by choosing $V_{\mu}$ in a convenient way.
In order to get a hint on the structure of $V_{\mu}$, it is very enlightening to calculate the equal-time pb of the field and its (generalized) conjugate momentum:
$\{A_{\mu}(x_{0},\mathbf{x}),\theta_{\lambda\nu}(x_{0},\mathbf{x}^{\prime})\}=\frac{1}{4}\left[\frac{\delta^{2}(0)}{8\pi^{3}}\right]\eta_{\mu\nu}K_{\lambda}(\mathbf{x},\mathbf{x}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part},$ (17)
where we have defined $K_{\mu}$ as
$K_{\mu}(\mathbf{x},\mathbf{x}^{\prime})=\frac{1}{4\pi^{3}}\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})k_{0}^{-1}\left[V_{\mu}\sin(\mathbf{k}\cdot\mathbf{x})\sin(\mathbf{k}\cdot\mathbf{x}^{\prime})+\frac{k_{\alpha}V^{\alpha}}{k_{\alpha}k^{\alpha}}k_{\mu}\cos(\mathbf{k}\cdot\mathbf{x})\cos(\mathbf{k}\cdot\mathbf{x}^{\prime})\right],$ (18)
since this result is independent of the auxiliary variables used to describe the system. Thus, Eq. (17) calls for taking $V_{\mu}$ equal to $ak_{0}k_{\mu}$, with $a$ a proportionality constant, so that
$\{A_{\mu}(x_{0},\mathbf{x}),\theta_{\lambda\nu}(x_{0},\mathbf{x}^{\prime})\}=\frac{a}{4}\left[\frac{\delta^{2}(0)}{8\pi^{3}}\right]\eta_{0\lambda}\eta_{\mu\nu}\delta^{(3)}(\mathbf{x}-\mathbf{x}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part}.$ (19)
With this selection, we easily identify the well-known result from standard field theory for the mentioned pb in the part of the latter equation that is associated with the free propagation of the electromagnetic field.
The generalization of the pb in Eq. (15) to the case of different times is immediate. Following Nikolić’s proposal [20], we establish the non-equal-time pb of functions $A$ and $B$ with respect to the canonical variables of the field as
$\{A,B\}({s})=a\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})k_{0}k_{\mu}\left(\frac{\partial{A}_{s}}{\partial{q}^{\nu}(s)}\frac{\partial{B}_{s}}{\partial\pi_{\mu\nu}(s)}-\frac{\partial{B}_{s}}{\partial{q}^{\nu}(s)}\frac{\partial{A}_{s}}{\partial\pi_{\mu\nu}(s)}\right),$ (20)
where $A_{s}$ and $B_{s}$ denote that functions $A$ and $B$ have been written in terms of the initial conditions $A(s)$ and $B(s)$, respectively, with $s$ some fixed instant. In the present case, and for the free field, we have that $A_{\mu}$ is a linear function of the initial conditions. Consequently, it is straightforward to calculate the non-equal-time pb for $A_{\mu}$ at two different space-time positions, so that we obtain that
$\{A_{\mu}(x_{a}),A_{\nu}(x_{a}^{\prime})\}({s})=a\pi{c}\left[\frac{\delta^{2}(0)}{8\pi^{3}}\right]\eta_{\mu\nu}\Delta(x_{\alpha}-x_{\alpha}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part},$ (21)
where $\Delta$ denotes the Lorentz-invariant Pauli-Jordan function. Similarly, one can directly calculate the non-equal-time pb of the vector potential and its conjugate momentum:
$\{A_{\mu}(x_{a}),\theta_{\lambda\nu}(x_{a}^{\prime})\}({s})=\frac{a}{4}\left[\frac{\delta^{2}(0)}{8\pi^{3}}\right]\eta_{\mu\nu}\partial_{\lambda}\Delta(x_{\alpha}-x_{\alpha}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part}.$ (22)
Both results are independent of $s$ and coincide, except for a numerical factor [^5] The factor $\delta^{2}(0)/8\pi^{3}$ in Eqs. (17), (19), (21), and (22), or similar factors, will appear when calculating pbs with respect to the canonical variables of quantities consisting of sums (integrals) of the contributions of all the modes of oscillation of the field, and is due to the particular choice of the normalization factor for these quantities. This factor could be “eliminated” through a renormalization process, but instead of sweeping it under the rug, we have decided to leave it explicit., with their counterparts from the conventional formulations of field theories (see, for example, Heitler [21]).
One more comment on Eqs. (21) and (22) is worth making as they show that the proposal in Eq. (20) leads to manifestly Lorentz-covariant results. Thus, although the proportionality of $V_{\mu}$ to $k_{0}$ might seem to break the covariance of the expression, we have actually that such proportionality transforms the pb with respect to variables $q_{\mu}$ and $\pi_{\mu\nu}$ into an effectively Lorentz-covariant form.
4 Reduction of the Proposal to the Standard (Non-covariant) Form of the Poisson Bracket
With the aim of finding an appropriate value for the proportionality constant $a$, let us explore the reduction process of the pb in Eq. (15) to its non-covariant form. First, and for purposes that will be clarified later, we rewrite Eq. (15) in terms of amplitudes $\mathcal{A}_{\mu}$ and their complex conjugates, so that we establish that
$\{A,B\}_{q\pi}=4\pi{i}c\{A,B\}_{\mathcal{AA}^{\ast}},$ (23)
where we have introduced the subscript $q\pi$ or $\mathcal{AA}^{\ast}$ to denote the set of variables with which the pb is computed and defined $\{A,B\}_{\mathcal{AA}^{\ast}}$ as
$\{A,B\}_{\mathcal{AA}^{\ast}}=a\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})k_{0}^{2}\left(\frac{\partial{A}}{\partial\mathcal{A}^{\mu}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{\mu}}-\frac{\partial{B}}{\partial\mathcal{A}^{\mu}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{\mu}}\right).$ (24)
Naturally, Eq. (24) does not correspond to the pb associated with the normal variables of the electromagnetic field, since for each oscillation mode of the field there is a sum of four terms instead of two, as occurs in the non-covariant formulation.
Nevertheless, we can proceed by assuming that there exists an inertial frame of reference such that we can write each amplitude $\mathcal{A}_{\mu}$ as
$\mathcal{A}_{\mu}=\mathcal{A}_{0}\epsilon_{\mu}^{0}+\mathcal{A}_{1}\epsilon_{\mu}^{1}+\mathcal{A}_{2}\epsilon_{\mu}^{2}+\mathcal{A}_{3}\epsilon_{\mu}^{3},$ (25)
i.e., as a sum of two transversal components, one longitudinal, and one scalar or temporal. In Eq. (25), vectors $\epsilon_{\mu}^{i}$, with $i=1,2,3$, are of the form $(0,\mathbf{e}_{i})$ where $\mathbf{e}_{1}$ and $\mathbf{e}_{2}$ denote polarization unit three-dimensional vectors that are orthogonal to each other and orthogonal to $\mathbf{e}_{3}$, a unit three-dimensional vector that is collinear to the spatial component of the wave vector of the mode under consideration. On the other hand, $\epsilon_{\mu}^{0}$ denotes a unit vector of the form $(1,\mathbf{0})$. Hence, we have that Eq. (24) can be written as
$\{A,B\}_{\mathcal{AA}^{\ast}}=a\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})k_{0}^{2}\left(\sum_{i=1}^{3}\left[\frac{\partial{A}}{\partial\mathcal{A}_{i}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{i}}-\frac{\partial{B}}{\partial\mathcal{A}_{i}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{i}}\right]-\left[\frac{\partial{A}}{\partial\mathcal{A}_{0}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{0}}-\frac{\partial{B}}{\partial\mathcal{A}_{0}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{0}}\right]\right).$ (26)
Then, since the Lorentz auxiliary condition, $\partial_{\mu}A^{\mu}=0$, implies that
$\mathcal{A}_{0}-\mathcal{A}_{3}=0,$ (27)
we have that the contributions of the scalar modes to the pb are compensated by those of the longitudinal modes for each oscillation mode of the field. Therefore, we find through this reduction process (Gupta-Bleuler process [22, 23]) that the pb in Eq. (26) is equal to
$\{A,B\}_{\mathcal{AA}^{\ast}}=\frac{a}{2}\sum_{\lambda=1}^{2}\int{d}^{3}k\hskip 1.42262ptk_{0}\left(\frac{\partial{A}}{\partial\mathcal{A}_{\lambda}}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{\lambda}}-\frac{\partial{B}}{\partial\mathcal{A}_{\lambda}}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{\lambda}}\right),$ (28)
with $k_{0}=|\mathbf{k}|$.
Now, the pb in Eq. (28) may not look like the standard pb, but it is in fact equal to the latter. To show this equivalence between expressions, we need to remember that the amplitudes in Eq. (28) are still four-dimensional and that the relation of these amplitudes with their three-dimensional counterparts is precisely [^6] See, for example, Heitler [21] or Bogoliubov and Shirkov [24].
$\mathcal{A}_{\mu}(\mathbf{k})=\frac{\mathcal{A}_{\mu}(k_{\alpha})}{\sqrt{2k_{0}}},$ (29)
so that we finally obtain
$\{A,B\}_{\mathcal{AA}^{\ast}}=\frac{a}{4}\sum_{\lambda=1}^{2}\int{d}^{3}k\left(\frac{\partial{A}}{\partial\mathcal{A}_{\lambda}(\mathbf{k})}\frac{\partial{B}}{\partial\mathcal{A}^{\ast}_{\lambda}(\mathbf{k})}-\frac{\partial{B}}{\partial\mathcal{A}_{\lambda}(\mathbf{k})}\frac{\partial{A}}{\partial\mathcal{A}^{\ast}_{\lambda}(\mathbf{k})}\right),$ (30)
which is the pb associated with the normal variables of the electromagnetic field if we take $a=4$ and identify $\lambda$ as the polarization index.
It is very important to note that although our proposal converges correctly to the standard pb, we have, due to Eq. (29), that
$\{\mathcal{A}_{\mu}(k_{\alpha}),\mathcal{A}^{\ast}_{\nu}(k_{\alpha}^{\prime})\}_{\mathcal{AA}^{\ast}}\neq\{\mathcal{A}_{\mu}(\mathbf{k}),\mathcal{A}^{\ast}_{\nu}(\mathbf{k}^{\prime})\}_{\mathcal{AA}^{\ast}},$ (31)
with
$\{\mathcal{A}_{\mu}(k_{\alpha}),\mathcal{A}^{\ast}_{\nu}(k_{\alpha}^{\prime})\}_{\mathcal{AA}^{\ast}}=4\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})k_{0}^{2}\eta_{\mu\nu}\delta^{(4)}(k_{\alpha}-k_{\alpha}^{\prime}).$ (32)
The pb in the right-hand side of Eq. (31) is generally used as the starting point for canonical quantization in the conventional formulations of quantum field theories. This primacy is the reason why the reduction process of the proposed pb was done in terms of amplitudes rather than canonical variables. It is particularly interesting to note that our result for the pb of two amplitudes of the electromagnetic field, Eq. (32), coincides in structure with that obtained by Dirac through a completely different approach [25].
5 Joint Description of the Electrodynamics of a Point Particle
As stated in Ref. [18], given the resemblance of the result of Eq. (16) with the canonical pb of classical mechanics, one can envision the possibility of defining a single pb associated with the canonical variables of the complete system. In order to construct such an object, let us first examine how to construct a Lorentz-covariant pb with respect to the canonical variables of the particle. One can find in the literature that this problem has been addressed on multiple occasions for purposes similar to those explored here (for example, see Ref. [25, 26, 27] and references therein). In any case, it is straightforward to propose the bilinear form
$\{A,B\}_{xp}=\frac{\partial{A}}{\partial{x}_{\mu}}\frac{\partial{B}}{\partial{p}^{\mu}}-\frac{\partial{A}}{\partial{x}_{\mu}}\frac{\partial{B}}{\partial{p}^{\mu}},$ (33)
as the equal-time pb of $A$ and $B$ with respect to the canonical variables of the particle (the generalization to the case of different times is direct from Nikolić’s proposal). Consequently, we have that the equal-time pb of the canonical variables of the mechanical subsystem is given by
$\{x_{\mu},p_{\nu}\}_{xp}=\eta_{\mu\nu}.$ (34)
On this basis, we establish that the set $(Q;P)$ of canonical variables of the particle+electromagnetic field system should be the union of the sets of canonical variables of the mechanical subsystem and the field, that is,
$(Q;P)=(x_{\mu},q_{\mu}(k_{\alpha});p_{\mu},\pi_{\mu\nu}(k_{\alpha})).$ (35)
Hence, we define
$\{A,B\}_{QP}(s)=\{A,B\}_{xp}(s)+\{A,B\}_{q\pi}(s)$ (36)
as the pb associated with all the canonical variables at time $s$ and postulate such bracket as the basis for a joint description of the whole system that is both canonical and Lorentz-covariant.
An interesting feature of the proposed description comes from the fact that the canonical variables of the system at any two different times, namely $s$ and $\tau>s$, are related through a canonical transformation. Thus, we have that
$\{x_{\mu}(\tau),p_{\nu}(\tau)\}_{QP}(s)=\{x_{\mu}(\tau),p_{\nu}(\tau)\}_{QP}(\tau),$ (37)
since pbs are invariant under those transformations. Therefore, we straightforwardly establish from Eqs. (34) and (37) that
$\{x_{\mu}(\tau),p_{\nu}(\tau)\}_{xp}(s)+\{x_{\mu}(\tau),p_{\nu}(\tau)\}_{q\pi}(s)=\eta_{\mu\nu},$ (38)
i.e, that the dependence of the canonical variables of the mechanical subsystem at time $\tau$ (or at any other time) on the initial conditions $x_{\mu}(s)$, $p_{\mu}(s)$, $q_{\mu}(s)$, and $\pi_{\mu\nu}(s)$ is precisely of the form necessary to ensure that their pb equals the metric tensor.
6 Conclusions
The apparent incompatibility between the requirement of field theories for locally describing the interactions between particles and the extended use of the canonical quantization process seems to have found, at least in principle, a solution in the mhf. As we showed in this article, it is feasible to use this formalism to construct a description of the complete (particle++electromagnetic field) system that is suitable for a canonical treatment and, at the same time, treats all space-time coordinates symmetrically. The pivotal element of such a description is an unambiguous Lorentz-covariant pb, Eqs. (15) and (20), which not only reproduces the well-known results from the conventional formulations of field theories, but also converges to the standard pb when reduced to its non-covariant form. On this basis, we obtained the pb of the canonical variables of the particle and of the field, Eqs. (16) and (34). These results constitute a generalization of the canonical pb of non-relativistic classical mechanics and might be taken as the departure point for the future development of a manifestly Lorentz-covariant quantization process.